OpenStudy (anonymous):

This prism has a volume of 90 cm3. What would the volume of the prism be if each dimension was tripled? (Scale factor is 3.) A. 2700 cm3 B. 270 cm3 C. 2430 cm3 D. 810 cm3

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

You can solve this 2 different ways.

OpenStudy (mathstudent55):

I will explain both ways to you.

OpenStudy (cwrw238):

if each dimension is * 3 then volume is * 3^3

OpenStudy (anonymous):

wait the answer is 270 right

OpenStudy (mathstudent55):

We saw earlier (the microwave oven problem) that the volume of a rectangular prism is: V = LWH

OpenStudy (anonymous):

is it 270

OpenStudy (mathstudent55):

We can check this problem. The dimensions are 6 cm, 5 cm, 3 cm. We multiply them all together, and we get V = 6 cm * 5 cm * 3 cm = 90 cm^3 That is what the problem tells us the original volume is.

OpenStudy (anonymous):

k

OpenStudy (mathstudent55):

The answer is not 270. We'll get there soon.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

The problem tells us each measure is multiplied by 3. Now let's multiply each measure by 3 and find the volume again.

OpenStudy (anonymous):

k

OpenStudy (mathstudent55):

6 cm * 3 = 18 cm 5 cm * 3 = 15 cm 3 cm * 3 = 9 cm Now we find the volume using our new dimensions that are 3 times larger than the original dimensions: V = 18 cm * 15 cm * 9 cm =

OpenStudy (mathstudent55):

What do you get?

OpenStudy (anonymous):

all of it i think the answer is 2430

OpenStudy (mathstudent55):

Correct. Remember it's with cm^3 units.

OpenStudy (mathstudent55):

That was the first way of solving the problem. Here is the second way.

OpenStudy (anonymous):

k here is a medal im going to post up another question

OpenStudy (mathstudent55):

If you have a solid and you apply a scale factor to its side of x, the area increases by a factor of x^2,and the volume increases by a factor of x^3.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Here, the scale factor on the side is 3, so the scale factor on the volume is 3^3. 3^3 = 3 * 3 * 3 = 27 That means the new volume is 27 times larger than the original one. 90 cm^3 * 27 = 2430 cm^3